OpenStudy (anonymous):

The function y=x^4 + bx^2 + 8x + 1 has a horizontal tangent and a point of inflection for the same value of x. What must be the value of b?

myininaya (myininaya):

is b the y-intercept?

myininaya (myininaya):

of the tangent line?

myininaya (myininaya):

oh i see

myininaya (myininaya):

nvm i see the b lol

myininaya (myininaya):

set y'=y''

OpenStudy (anonymous):

Okay!

myininaya (myininaya):

i dont know if that will help us lets see what heppens

myininaya (myininaya):

oh wait erase that we have 2 equations right the one with y' and the one with y'' set both = to zero so y'' gives us 12x^2+2b=0 which means b=-12x^2/2=-6x^2 so take y' which gives us 4x^2+2bx+8=0 and where you see that x plug in -6x^2

myininaya (myininaya):

so we have 4x^3+2(-6x^2)x+8=0 4x^3-12x^3+8=0 -8x^3+8=0 and so on...

myininaya (myininaya):

you did see the typo above right when i wrote 4x^2 its suppose to be 4x^3

OpenStudy (anonymous):

Yea lol! I'm going to see what I get!

myininaya (myininaya):

so anyways we have 8x^3=8 so x^3=1 so x=1

myininaya (myininaya):

if x=1 and b=-6x^2, then what does b=__

OpenStudy (anonymous):

b=-6 and that's what the answer says!! :D

myininaya (myininaya):

did you understand the steps? do you want me to say why i did something or do you got it? this was a nice problem

OpenStudy (anonymous):

I understand! Yea it was a nice problem indeed!

myininaya (myininaya):

i seen another typo way up there it says replace c with -6x^2 thats suppose to be replace b with -6x^2 i cant type im sorry

myininaya (myininaya):

not c but x lol

OpenStudy (anonymous):

You were probably typing fast but yea I understand! Thanks!

myininaya (myininaya):

ok have funs!